The rate mass of a fluid is 12 grams/s, is flowing in a tube at 0.9 m/s and it has a density of 1.7 grams/m3. What is the diameter of the tube?


1
Expert's answer
2021-03-04T11:51:51-0500
"\frac{dm}{dt}=Av\rho \\\frac{dm}{dt}=0.25\pi d^2v\rho 
\\12=0.25\pi d^2(0.9)(1.7)\\d=3.16\ m"